x = 4y + 5
x - 3y = -2

Answer by stanbon(75887)

x = 4y + 5
x - 3y = -2
--------------------
Substitute for "x" and solve for "y":
4y+5 - 3y = -2
y + 5 = -2
y = -7
===========
Since x = 4y+5 ; x = 4*-7+5 ; x = -23
=====================
Cheers,
Stan H.